OpenStudy (anonymous):

an investor invested a total of 2100 in two mutual funds one fund earned 4% while the other earned a 2% profit if the investors total profit was 70 how much was invested into each fund

OpenStudy (anonymous):

please help!!!

OpenStudy (nurali):

Total investment = $2100 Amount invested for 4% profit = x Amount invested for 2% profit = 2100-x Total Profit = $70 (Amount invested for 4% profit)(4%)+ (Amount invested for 2% profit)(2%)= Total Profit (x)(4/100)+ (2100-x)(2/100)= 70 Multiply by 100 both sides (100)(x)(4/100)+ (100)(2100-x)(2/100)= 70*100 4x+(2100-x)(2)= 7000 4x+4200-2x=7000 2x=7000-4200 2x=2800 x=2800/2 x=1400 Amount invested for 4% profit = x = 1400 Amount invested for 2% profit = 2100-x = 2100-1400=700 Check: 4% of 1400 = (4/100)(1400)=56 2% of 1100 = (2/100)(700)=14 Total profit = 56+14=70